# Complete transparency with three active-passive-coupled optical resonators

### v1

@@ Abstract
The phenomena of induced transparency, with the typical examples of electromagnetically induced transparency (EIT) in atomic media and coupled optical resonators, have attracted tremendous interest since their discoveries. Due to the limitations of the involved elements, however, near-100% transmissions were reported under highly demanding experimental conditions for atomic and other media. Based on a structure of three linearly coupled optical resonators, an active one carrying a possibly arbitrary optical gain and two passive ones simply with dissipation, we demonstrate that a transmitted light field can become completely transparent through the structure, which displays all properties similar to those of EIT. Manifested by a destructive interference to annihilate the intracavity field in the resonator directly coupled to the input, this complete transparency exists for any feasible power of the transmitted field and all realizable coupling strengths of the dark resonator with the input and the neighboring resonator, as long as the inter-cavity coupling for two other resonators is adjustable over a suitable range. A free control on the transparency window size and output field intensity can be realized by tuning two inter-cavity couplings without modifying the built-in system parameters.
@@
Since its first observation with a three-level atomic medium [1, 2], the phenomenon of EIT has been extensively studied together with the accompanying slow-light effect [3, 4, 5, 6, 8, 7, 9]. Due to non-zero decay rates of the involved atomic levels and additional absorption pathways in the presence of hyperfine or Zeeman sublevels, achieving a complete transparency under EIT condition is difficult. So far, near-100100% transmissions were reported with very few setups, such as optically dense rare-earth doped crystal [4], atomic vapor excited with circularly polarized laser of multimode [6], and single artificial atoms of superconductor [10]. On the other hand, there exist the EIT-like phenomena in coupled optical resonators known as coupled-resonator-induced transparency (CRIT) [11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and they have developed into a prominent research direction [21, 22, 23]. Two linearly coupled dissipative resonators display all properties similar to those of three-level atomic EIT. The addition of nonlinear optical elements into optical resonator also leads to various types of induced transparency [23, 24, 26, 25, 27, 28, 29]. Especially, cavity optomechanical systems were found to demonstrate optomechanically induced transparency (OMIT) [30, 31, 32], which has stimulated tremendous interests in its research [33, 34, 35, 38, 36, 37, 39, 40, 41, 42]. To these cavity systems, it is equally challenging to realize a complete transparency, due to the unavoidable dissipation of the involved elements such as optical resonator or mechanical resonator with a finite quality factor.
@@
Here, we present an experimentally feasible approach to realizing the complete transparency with three linearly coupled resonators [see Fig. 1(a)]. The structure consists of one active resonator with optical gain and two other dissipative ones. A special cavity dark mode with totally vanishing intracavity field in the dissipative resonator directly coupled to the input emerges under certain conditions, and it realizes the 100% transmission at the center of a transparency window. Unlike various schemes relying on field amplification mechanisms [8, 20, 28, 43], the complete intracavity field cancellation is realizable with an arbitrary optical gain in the neighboring resonator, no matter how large the field loss in the dark cavity can be. Contrary to intuition, it is not necessary to have an correspondingly enhanced gain to balance any amount of increased field damping for the dark cavity, as long as the inter-cavity coupling between two other resonators can be adjusted over a suitable range. Different from atomic EIT and many other induced-transparency scenarios such as CRIT and OMIT, this phenomenon is based on a previously unknown destructive interference that is irrespective of the pump field intensity and the coupling strengths of the others to the dark resonator.
@@
For the structure shown in Fig. 1(a), the symbols $\hat{a}_{i}$ and $\omega_{i}$ stand for the intracavity mode and the resonance frequency of cavity $i$ ($i=1,2,3$), respectively. The pump field transmitted at the frequency $\omega_{p}$ and with the amplitude $\varepsilon_{p}=\sqrt{P/(\hbar\omega_{p})}$ ($P$ is its power) couples to cavity $1$. The active resonator, cavity $2$, has the net gain rate $\kappa_{2}=\kappa_{2,0}/\left(1+\frac{<\hat{a}^{\dagger}_{2}\hat{a}_{2}>}{I_{S}}\right)-\gamma_{2}>0$ ($I_{S}$ is the gain saturation intensity and $\gamma_{2}$ the damping rate of the cavity) and is coupled to two other dissipative ones with their damping rates $\kappa_{1}$ and $\kappa_{3}$, respectively, at the inter-cavity couplings $J_{1}$ and $J_{2}$, which are determined by the distances between the resonators. Hence, similar to those of parity-time (PT) symmetric optical structures [47, 48, 49, 50, 51, 52], the effective Hamiltonian of the setup reads as ($\hbar=1$)
@@
where the damping rate $\kappa_{1}=\kappa_{ex}+\kappa_{in}$ of cavity $1$ comes from both external coupling with the pump and its intrinsic loss.
Figure 1: (a) The schematic structure of three linearly coupled microresonators. Resonator 22 carries a gain medium without showing the external pump used for the field amplification. The actual position of the optical fiber is within the plane of the microresonators. There will be a totally vanishing intracavity field $\hat{a}_{1}$, once resonator 22 and 33 satisfy the condition in Eq. (7). Then, the transmitted field will have the 100100% transmission at the center of a narrow frequency window, as shown by the example in the inset, which is obtained with $\kappa_{1}=2\kappa_{ex}=10\,\mathrm{MHz}$, $\kappa_{2}=0.2\,\mathrm{MHz}$, $\kappa_{3}=5\,\mathrm{MHz}$, and $J_{1}=2\,\mathrm{MHz}$, and $J_{2}=1\,\mathrm{MHz}$ for three identical microresonators. (b) An example of the evolving intracavity fields after including the optical gain saturation effect. The parameters are $\varepsilon_{p}=10^{4}\,\sqrt{\mathrm{MHz}}$ (equivalent to $P=12.8\,\mathrm{\mu W}$ at $\lambda=1550$ nm), $\kappa_{1}=2\kappa_{ex}=10\,\mathrm{MHz}$, $\kappa_{2,0}=0.2\,\mathrm{MHz}$, $I_{S}=10^{8}$, $\kappa_{3}=5\,\mathrm{MHz}$, $J_{1}=20\,\mathrm{MHz}$, and $J_{2}=1\,\mathrm{MHz}$. The saturated gain rate after the stabilization is $\kappa_{2}=0.195\,\mathrm{MHz}$, and the cavity damping rates are chosen to be on the levels of those in some recent experiments [44, 45, 46].
To capture the main physical picture, we will consider that the gain rate $\kappa_{2}$ quickly saturates to a constant value during the system evolution to a stability, so that the gain rate $\kappa_{2}$ can be treated as a constant. In addition to the previous studies on the relevant effects (see e.g., Refs. [53, 54, 55, 56, 57]), the properties of the gain saturation with a real-time $\kappa_{2}(t)$ in the current scenario, together with the associate gain noise effect [58, 59] which can be neglected in experiments [60], are discussed in Appendix A. Then the effective Hamiltonian in Eq. (1) leads to the following dynamical equation of the operator mean values $a_{i}(t)=\langle\hat{a}_{i}(t)\rangle$:
@@
In a stable regime (the real parts of all eigenvalues of M are not positive), the solution to Eq. (2) takes the form $a_{k}=A_{k}e^{-i\omega_{p}t}$ ($k=1,2,3$). Substituting this form into Eq. (2), we obtain the exact solution in the stability:
@@
with $\Delta_{i}=\omega_{p}-\omega_{i}$ ($i=1,2,3$) and $\eta=J_{2}^{2}(\kappa_{1}-i\Delta_{1})$. It can be seen from Eq. (4) that $A_{2}$ and $A_{3}$ cannot be zero because $J_{1}$, $J_{2}$ and $\kappa_{3}$ are non-zero, but $A_{1}$ can be zero under certain conditions. These conditions are crucial for achieving a complete transparency. In order to obtain the conditions, we rewrite $A_{1}$ in the form of continued fraction:
@@
Such continued fraction form can provide a deeper insight into the physical mechanism and helps to conveniently find the complete transparency conditions for the structures of more than three resonators (see Appendix B). The properties of a transmitted light field is determined by the induced $A_{1}$, with the real and imaginary part of the correspondingly defined susceptibility $\varepsilon_{T}=\sqrt{2\kappa_{ex}}A_{1}/\varepsilon_{p}$ (reformulated from a similar definition [30] according to our form of dynamical equation) respectively describing the absorptive and dispersive behavior of a single-frequency input to the structure.
@@
where $x=\omega_{p}-\omega_{0}$. In comparison, the susceptibility of two coupled passive resonators [18, 22] takes the form $\varepsilon_{T}=2\kappa_{ex}/[\gamma_{1}-ix+J^{2}/(\gamma_{2}-ix)]$, with the corresponding definitions for the coupling and damping rates. The denominator $\gamma_{2}-ix$ of its subfraction cannot be zero, due to the second cavity’s damping rate $\gamma_{2}\neq 0$ in reality.
Figure 2: (a) The distribution of the absolute susceptibility, $|\varepsilon_{T}|=\sqrt{2\kappa_{ex}}|A_{1}|/\varepsilon_{p}$, at $x=0$, with respect to two parameters $\sqrt{\kappa_{2}\kappa_{3}}$ and $J_{2}$. Here, we fix the rest parameters at $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{3}=5\,\mathrm{MHz}$, and $J_{1}=10\,\mathrm{MHz}$. The calculation is based on Eq. (5) but with three identical resonance frequencies. (b) The distribution of $\text{Re}(\varepsilon_{T})$ with respect to $x$ and $J_{1}$. Here, the rest of the parameters are set to be $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{2}=0.05\,\mathrm{MHz}$, $\kappa_{3}=20\,\mathrm{MHz}$, and $J_{2}=1\,\mathrm{MHz}$. The calculation is based on Eq. (9).
The solution of the steady intracavity fields,
$(A_{1},A_{2},A_{3})^{\mathrm{T}}=\sqrt{2\kappa_{ex}}\varepsilon_{p}\left(0,\frac{i}{J_{1}},-\sqrt{\frac{\kappa_{2}}{\kappa_{3}}}\frac{1}{J_{1}}\right)^{\mathrm{T}},$ (10)
under the conditions of complete transparency in Eq. (8), can be proved to be the zero eigenstate of the effective Hamiltonian in Eq. (1) with $\kappa_{2}=\text{constant}$; see Appendix C. Similar to the definition for the networks of linearly coupled bosonic modes [61], the mode $\hat{a}_{1}$ with $A_{1}=0$, which is decoupled from all possible excitations of $\hat{a}_{2}$ and $\hat{a}_{3}$, is regarded as a cavity dark mode. Different from those formed by atom-photon interactions [62, 63, 64] and optomechanical interaction [65, 66], this cavity dark mode is simply created by the linear couplings between the cavities, together with an optical gain in one of them. In this vacuum state $|0\rangle$ of cavity $1$, the pump action $\hat{H}_{1}=i\sqrt{2\kappa_{ex}}\varepsilon_{p}(\hat{a}_{1}^{\dagger}-\hat{a}_{1})$ and the coupling action $\hat{H}_{2}=-J_{1}(A_{2}\hat{a}_{1}^{\dagger}+A_{2}^{\ast}\hat{a}_{1})$ from cavity $2$ could affect its photon number (see Appendix C). However, one has the exact cancellation
@@
with $A_{2}=i\sqrt{2\kappa_{ex}}\varepsilon_{p}/J_{1}$ in Eq. (10), indicating an absolutely forbidden transition to the excited states of cavity $1$. This perfect cancellation exists for all possible values of the directly involved system parameters ($J_{1}$ and $\sqrt{2\kappa_{ex}}\varepsilon_{p}$), in contrast to the destructive interference mechanisms in the previously known scenarios of induced transparency. Tuning the coupling strength $J_{2}$ (with the examples on microspheres [15] and microtoroid resonators [67]) so that the system goes close to the diagonal line in Fig. 2(a), one will obtain an approximate dark mode in the red-colored region. Varying the other coupling $J_{1}$ will modify the distribution of the approximate dark mode with respect to the detuning $x$, as shown in Fig. 2(b). It is seen from the continued fraction form in Eq. (5) that no matter how small $J_{1}$ is, there exist the approximate dark modes in the neighborhood of the parameter submanifolds determined by Eq. (6).
Once the mode $\hat{a}_{1}$ is tuned into a dark mode, the fields in cavity $2$ and $3$ will exchange the photons with each other in a self-sustained way, i.e., their field losses happen to be compensated by the optical gain, having their photon numbers at the fixed ratio $\langle\hat{a}_{2}^{\dagger}\hat{a}_{2}\rangle/\langle\hat{a}_{3}^{\dagger}\hat{a}_{3}\rangle=\kappa_{3}/\kappa_{2}$ without any supply from the pump field. This counter-intuitive phenomenon occurs with any gain rate $\kappa_{2}$, as long as the coupling $J_{2}$ can be correspondingly adjusted according to Eq. (8). At the limit $\kappa_{2}\ll\kappa_{3}$ another resonator, cavity $3$, will tend to a dark one as well.
The most important features of the structure are reflected in the absorptive curve of $\text{Re}(\varepsilon_{T})$ and the dispersive curve of $\text{Im}(\varepsilon_{T})$. In a situation of $\kappa_{2}<\kappa_{3}$, these curves are exemplified by those in Fig. 3(a1). The size of the transparency window, the full width at half maximum (FWHM) of the curve $\text{Re}(\varepsilon_{T})$ around $x=0$, can be approximated as $J_{1}^{2}/\kappa_{1}$ when $J_{1}$ is small; cf. the approximate transparency window width $|\Omega|^{2}/\gamma$ ($\Omega$ is the pump Rabi frequency and $\gamma$ the level decay rate) of three-level atomic EIT. Correspondingly, the slope of the curve of $\text{Im}(\varepsilon_{T})$ has the exact form,
@@
Figure 3: (a1)-(a2) The examples of $\text{Re}(\varepsilon_{T})$ and $\text{Im}(\varepsilon_{T})$, as well as the time delays of the transmitted fields, in the regime of $\kappa_{2}<\kappa_{3}$. We use $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{2}=0.01\,\mathrm{MHz}$, $\kappa_{3}=20\,\mathrm{MHz}$, $J_{1}=0.01\,\mathrm{MHz}$ in (a), and $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$, so that the transparency window size is in the order of $\mathrm{Hz}$ and the time delays approach the order of second. (b1)-(b2) The other examples in the regime of $\kappa_{2}>\kappa_{3}$. The system parameters are set to be $\kappa_{1}=2\kappa_{ex}=10\,\mathrm{MHz}$, $\kappa_{2}=1\,\mathrm{MHz}$, $\kappa_{3}=0.1\,\mathrm{MHz}$, and $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$. In (b1) we have $J_{1}=4\,\mathrm{MHz}$.
@@
Figure 4: (a) The comparison between the output/input ratios due to three different couplings $J_{2}$, which are, respective smaller than (the red curve), equal to (the blue curve), and larger than (the black curve) the value $J_{2,c}$ of the $100\%$ transmission at the transparency window center. Represented by these three samples, the output can be continuously amplified or reduced by tuning the coupling $J_{2}$ across the point $J_{2,c}$, while the built-in optical gain is unchanged. (b) The corresponding dispersion curves. Here, the fixed system parameter are chosen as $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{2}=0.06\,\mathrm{MHz}$, $\kappa_{3}=6\,\mathrm{MHz}$, and $J_{1}=1\,\mathrm{MHz}$.
@@
Instead of being a limit point that should be asymptotically approached, the $100\%$ transmission rate emerges at a midpoint of tuning the coupling $J_{2}$. As represented by the sample results in Fig. 4(a), tuning the coupling strength from $J_{2}>J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ to $J_{2}<J_{2,c}$ outputs a light field continuously varying from the smaller than the $100\%$ transmission to the larger than the $100\%$ transmission at the transparency window center. The dispersion property demonstrating the slow-light effect is still preserved in the regimes deviating from the complete intracavity field cancellation; see Fig. 4(b). As the distance between cavity $2$ and $3$ keeps increasing from where cavity $1$ is in the vacuum state of completely eliminated intracavity field, which is at $J_{2,c}=0.6$ MHz for the system described in Fig. 4, more generated photons will go to cavity $1$ so that a higher and higher ratio $|a_{out}/a_{in}|^{2}>1$ will be achieved until the system loses its dynamical stability at $J_{2}=0.2449$ MHz. At a very small coupling $J_{2}$, this system under the condition $J_{1}<\sqrt{\kappa_{1}\kappa_{2}}$ behaves like two active-passive-coupled resonators in the PT-symmetry broken regime. Although these phenomena of amplifying and reducing the output under a fixed gain rate $\kappa_{2}$ are illustrated with the parameters of high-quality microresonators [44, 45, 46], there is no restriction on the qualities of the used optical resonators, since the conditions on system stability given in Appendix D and a tunable range covering the point $J_{2,c}$ guarantee their occurrence.

### v2

@@ Abstract
The phenomena of induced transparency, with the typical examples of electromagnetically induced transparency (EIT) in atomic media and coupled optical resonators, have attracted tremendous interest since their discoveries. Due to the limitations of the involved elements, however, near-100% transmissions were reported under highly demanding experimental conditions for atomic and other media. Based on a structure of three linearly coupled optical resonators, an active one carrying a possibly arbitrary optical gain and two passive ones simply with dissipation, we demonstrate that a transmitted light field can become completely transparent through the structure, which displays all properties similar to those of EIT. A destructive interference mechanism is discovered to totally eliminate the intracavity field in the resonator directly coupled to the transmitted field of any feasible power, when the coupling strength of two other resonators is tuned to a point inside the designed range. This mechanism works for all possible coupling strengths of the dark resonator with its neighboring resonator and the input field, as well as for any available quality factor from its fabrication. The transparency window size and the output field intensity at the transparency window center can be freely engineered by adjusting two inter-cavity couplings, without modifying the built-in optical gain which can be just slightly stronger than the dissipation of the active resonator.
@@
Since its first observation with a three-level atomic medium [1, 2], the phenomenon of EIT has been extensively studied together with the accompanying slow-light effect [3, 4, 5, 6, 8, 7, 9]. Due to non-zero decay rates of the involved atomic levels and additional absorption pathways in the presence of hyperfine or Zeeman sublevels, achieving a complete transparency under EIT condition is difficult. So far, near-100100% transmissions were reported with very few setups, such as optically dense rare-earth doped crystal [4], atomic vapor excited with circularly polarized laser of multimode [6], and single artificial atoms of superconductor [10]. On the other hand, there exist the EIT-like phenomena in coupled optical resonators known as coupled-resonator-induced transparency (CRIT) [11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and they have developed into a prominent research direction [21, 22, 23]. Two linearly coupled dissipative resonators display all properties similar to those of three-level atomic EIT. The addition of nonlinear elements into optical resonator also leads to various types of induced transparency [23, 24, 26, 25, 27, 28, 29]. Especially, cavity optomechanical systems were found to demonstrate optomechanically induced transparency (OMIT) [30, 31, 32], which has stimulated tremendous interests in its research [33, 34, 35, 38, 36, 37, 39, 40, 41, 42]. To these cavity systems, it is equally challenging to realize a complete transparency, due to the unavoidable dissipation of the involved elements such as optical resonator or mechanical resonator with a finite quality factor.
@@
Here, we present an experimentally feasible approach to realizing the complete transparency with three linearly coupled resonators [see Fig. 1(a)]. The structure consists of one active resonator with optical gain and two other dissipative ones. A special cavity dark mode with totally vanishing intracavity field in the dissipative resonator directly coupled to the input emerges under certain conditions, and it realizes the 100% transmission at the center of transparency window. Unlike various schemes [8, 20, 43] where optical gain is used to compensate for optical loss, the complete intracavity field cancellation is realizable with an arbitrary net gain in the neighboring resonator, no matter how large the field loss in the dark cavity can be. Contrary to intuition, it is not necessary to have an correspondingly enhanced gain to balance any amount of increased field damping for the dark cavity, as long as the inter-cavity coupling between two other resonators can be adjusted over a suitable range. Different from atomic EIT and many other induced-transparency scenarios such as CRIT and OMIT, this phenomenon is based on a previously unknown interference mechanism that is irrespective of the transmitted field intensity and the coupling strengths of the others to the dark resonator.
@@
For the structure shown in Fig. 1(a), the symbols $\hat{a}_{i}$ and $\omega_{i}$ stand for the intracavity mode and the resonance frequency of cavity $i$ ($i=1,2,3$), respectively. The transmitted field at the frequency $\omega_{p}$ and with the amplitude $\varepsilon_{p}=\sqrt{P/(\hbar\omega_{p})}$ ($P$ is its power) couples to cavity $1$. The active resonator, cavity $2$, has the net gain rate $\kappa_{2}=\kappa_{2,0}/\left(1+\frac{<\hat{a}^{\dagger}_{2}\hat{a}_{2}>}{I_{S}}\right)-\gamma_{2}>0$ ($\kappa_{2,0}$ is the initial gain rate, $I_{S}$ the gain saturation intensity, and $\gamma_{2}$ the cavity’s damping rate) and is coupled to two other dissipative ones with their damping rates $\kappa_{1}$ and $\kappa_{3}$, respectively, at the inter-cavity coupling strengths $J_{1}$ and $J_{2}$, which are determined by the distances between the resonators. Hence, similar to those of parity-time (PT) symmetric optical structures [47, 48, 49, 50, 51, 52], the effective Hamiltonian of the setup reads as ($\hbar=1$)
@@
where the damping rate $\kappa_{1}=\kappa_{ex}+\kappa_{in}$ of cavity $1$ comes from both external coupling and intrinsic loss.
Figure 1: (a) The schematic structure of three linearly coupled microresonators. Resonator 22 carries a gain medium without showing its associated pump. The actual position of the optical fiber is within the plane of the microresonators. There will be a totally vanishing intracavity field $\hat{a}_{1}=0$, once resonator 22, 33 and a single-frequency input satisfy the conditions in Eq. (7). Then, the transmitted fields will have the 100100% transmission at a transparency window center, as exemplified in the inset obtained with $\kappa_{1}=2\kappa_{ex}=10\,\mathrm{MHz}$, $\kappa_{2}=0.2\,\mathrm{MHz}$, $\kappa_{3}=5\,\mathrm{MHz}$, $J_{1}=2\,\mathrm{MHz}$, and $J_{2}=1\,\mathrm{MHz}$, for three microresonators of the same resonance frequency. (b) An example of the evolving intracavity fields under gain saturation effect. The parameters are $\varepsilon_{p}=10^{4}\,\sqrt{\mathrm{MHz}}$ (equivalent to $P=12.8\,\mathrm{\mu W}$ at $\lambda=1550$ nm), $\kappa_{1}=2\kappa_{ex}=10\,\mathrm{MHz}$, $\kappa_{2,0}=0.2\,\mathrm{MHz}$, $I_{S}=10^{8}$, $\kappa_{3}=5\,\mathrm{MHz}$, $J_{1}=20\,\mathrm{MHz}$, and $J_{2}=1\,\mathrm{MHz}$. The saturated gain rate after the stabilization is $\kappa_{2}=0.195\,\mathrm{MHz}$, and the cavity damping rates are chosen to be on the levels of those in some recent experiments [44, 45, 46].
To capture the main physical picture, we will consider a constant net gain rate $\kappa_{2}$ while the system is driven by a transmitted field. In addition to the previous studies (see e.g., Refs. [53, 54, 55, 56, 57]), the effect of gain saturation with a real-time $\kappa_{2}(t)$ during the system’s dynamical evolution [like the process in Fig. 1(b)], together with the associate gain noise effect [58, 59] which can be neglected in experiments [60], is discussed in Appendix A and D. Then the effective Hamiltonian in Eq. (1) leads to the following dynamical equation of the mean modes $a_{i}(t)=\langle\hat{a}_{i}(t)\rangle$ ($i=1,2,3$):
@@
In a stable regime (the real parts of all eigenvalues of M are not positive), the solution to Eq. (2) takes the form $a_{k}=A_{k}e^{-i\omega_{p}t}$ ($k=1,2,3$). Substituting this form into Eq. (2), we obtain the exact solution:
@@
with $\Delta_{i}=\omega_{p}-\omega_{i}$ ($i=1,2,3$) and $\eta=J_{2}^{2}(\kappa_{1}-i\Delta_{1})$. It can be seen from Eq. (4) that $A_{2}$ and $A_{3}$ cannot be zero because $J_{1}$, $J_{2}$ and $\kappa_{3}$ are non-zero, but $A_{1}$ can be zero under certain conditions. These facts are crucial for achieving a complete transparency. In order to obtain the conditions, we rewrite $A_{1}$ in the form of continued fraction:
@@
Such continued fraction forms help to find the complete transparency conditions for all similar structures (see Appendix B). The properties of a transmitted light field is determined by the induced $A_{1}$, with the real and imaginary part of the susceptibility $\varepsilon_{T}=\sqrt{2\kappa_{ex}}A_{1}/\varepsilon_{p}$ (reformulated from a similar definition [30] by our form of dynamical equation) respectively describing the absorptive and dispersive behavior of a single-frequency input to the structure.
@@
where $x=\omega_{p}-\omega_{0}$. In comparison, the susceptibility of two coupled passive resonators [18, 22, 23] takes the form $\varepsilon_{T}=2\kappa_{ex}/[\kappa_{1}-ix+J^{2}/(\gamma_{2}-ix)]$, with the corresponding definitions for the coupling and damping rates. The denominator $\gamma_{2}-ix$ of its subfraction cannot be zero, due to the second cavity’s damping rate $\gamma_{2}\neq 0$ in reality.
Figure 2: (a) The distribution of the absolute susceptibility, $|\varepsilon_{T}|=\sqrt{2\kappa_{ex}}|A_{1}|/\varepsilon_{p}$, at $x=0$, with respect to two parameters $\sqrt{\kappa_{2}\kappa_{3}}$ and $J_{2}$. Here, we fix the rest parameters at $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{3}=5\,\mathrm{MHz}$, and $J_{1}=10\,\mathrm{MHz}$. The calculation is based on Eq. (5) but with the same cavity resonance frequency. (b) The distribution of $\text{Re}(\varepsilon_{T})$ with respect to $x$ and $J_{1}$. Here, the rest of the parameters are set to be $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{2}=0.05\,\mathrm{MHz}$, $\kappa_{3}=20\,\mathrm{MHz}$, and $J_{2}=1\,\mathrm{MHz}$. The calculation is based on Eq. (9).
The stabilized intracavity fields,
$(A_{1},A_{2},A_{3})=\sqrt{2\kappa_{ex}}\varepsilon_{p}\left(0,\frac{i}{J_{1}},-\sqrt{\frac{\kappa_{2}}{\kappa_{3}}}\frac{1}{J_{1}}\right)$ (10)
under the conditions of complete transparency in Eq. (8), correspond to the zero eigenstate of the effective Hamiltonian in Eq. (1) (see the proof in Appendix C). Similar to the definition for a network of linearly coupled bosonic modes [61], the mode $\hat{a}_{1}$ with $A_{1}=0$, which is decoupled from all possible excitations of $\hat{a}_{2}$ and $\hat{a}_{3}$, is regarded as a cavity dark mode. Different from those formed by atom-photon interactions [62, 63, 64] and optomechanical interaction [65, 66], this cavity dark mode is simply created by the linear couplings between the cavities, together with an optical gain in one of them. In this vacuum state $|0\rangle$ of cavity $1$, the driving action $\hat{H}_{1}=i\sqrt{2\kappa_{ex}}\varepsilon_{p}(\hat{a}_{1}^{\dagger}-\hat{a}_{1})$ of the transmitted field and the coupling action $\hat{H}_{2}=-J_{1}(A_{2}\hat{a}_{1}^{\dagger}+A_{2}^{\ast}\hat{a}_{1})$ from cavity $2$ could affect its photon number (see Appendix C). However, one has the exact cancellation
@@
with the stabilized $A_{2}=i\sqrt{2\kappa_{ex}}\varepsilon_{p}/J_{1}$ in Eq. (10), indicating an absolutely forbidden transition to the excited states of cavity $1$. This perfect cancellation exists for all possible values of the directly involved system parameters ($J_{1}$ and $\sqrt{2\kappa_{ex}}\varepsilon_{p}$), in contrast to the destructive interference mechanisms in the previously known scenarios of induced transparency. Tuning the coupling strength $J_{2}$ (with the examples on microspheres [15], ring resonators [20], and microtoroid resonators [18, 67]) so that the system goes close to the diagonal line in Fig. 2(a), one will obtain an approximate vacuum state of cavity $1$. Varying the other coupling $J_{1}$ will modify the distribution of this approximate dark mode with respect to the detuning $x$; see Fig. 2(b). The continued fraction form in Eq. (5) shows that no matter how small $J_{1}$ is, there exist the approximate dark modes in the neighborhood of the parameter submanifolds determined by Eq. (6).
Once the mode $\hat{a}_{1}$ is tuned into a dark one, the fields in cavity $2$ and $3$ will exchange the photons with each other in a self-sustained way, having their photon numbers at the fixed ratio $\langle\hat{a}_{2}^{\dagger}\hat{a}_{2}\rangle/\langle\hat{a}_{3}^{\dagger}\hat{a}_{3}\rangle=\kappa_{3}/\kappa_{2}$ without any supply from the input field. The stabilized photon number $\langle\hat{a}_{2}^{\dagger}\hat{a}_{2}\rangle$ in the active resonator, which counter-intuitively becomes independent of its net gain rate $\kappa_{2}$ [see Eq. (10)], is unaffected by their coupling strength $J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ in this state. For any such system that can be tuned to this point $J_{2,c}$, the designed gain rate can be just slightly higher than the damping rate $\gamma_{2}$, to have a net gain rate $\kappa_{2}$ as small as possible with a correspondingly increased $\kappa_{3}$.
The most important features of the structure are reflected in the absorptive curve of $\text{Re}(\varepsilon_{T})$ and the dispersive curve of $\text{Im}(\varepsilon_{T})$. In a situation of $\kappa_{2}<\kappa_{3}$, these curves after the coupling strength $J_{2}$ is tuned to $J_{2,c}$ are exemplified by those in Fig. 3(a1). The size of the transparency window, the full width at half maximum (FWHM) of the curve $\text{Re}(\varepsilon_{T})$ around $x=0$, can be approximated as $J_{1}^{2}/\kappa_{1}$ when $J_{1}$ is small; cf. the approximate transparency window width $|\Omega|^{2}/\gamma$ ($\Omega$ is the pump Rabi frequency and $\gamma$ the level decay rate) of three-level atomic EIT. Correspondingly, the slope of the curve of $\text{Im}(\varepsilon_{T})$ has the exact form,
@@
Figure 3: (a1)-(a2) The examples of $\text{Re}(\varepsilon_{T})$ and $\text{Im}(\varepsilon_{T})$, as well as the time delays of the transmitted fields, in the regime of $\kappa_{2}<\kappa_{3}$. We use $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{2}=0.01\,\mathrm{MHz}$, $\kappa_{3}=20\,\mathrm{MHz}$, $J_{1}=0.01\,\mathrm{MHz}$ in (a1), and $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$, so that the transparency window size is in the order of $\mathrm{Hz}$ and the time delays approach the order of second. (b1)-(b2) The other examples in the regime of $\kappa_{2}>\kappa_{3}$. The system parameters are set to be $\kappa_{1}=2\kappa_{ex}=10\,\mathrm{MHz}$, $\kappa_{2}=1\,\mathrm{MHz}$, $\kappa_{3}=0.1\,\mathrm{MHz}$, and $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$. In (b1) we have $J_{1}=4\,\mathrm{MHz}$.
@@
Figure 4: (a) The comparison between the output/input ratios due to three different couplings $J_{2}$, which are, respective smaller than (the red curve), equal to (the blue curve), and larger than (the black curve) the value $J_{2,c}$ of the $100\%$ transmission at the transparency window center. Represented by these three samples, the output can be continuously amplified or reduced by tuning the coupling $J_{2}$ across the point $J_{2,c}$, while the built-in optical gain is unchanged. (b) The corresponding dispersive curves. Here, the fixed system parameter are chosen as $\kappa_{1}=2\kappa_{ex}=20\,\mathrm{MHz}$, $\kappa_{2}=0.06\,\mathrm{MHz}$, $\kappa_{3}=6\,\mathrm{MHz}$, and $J_{1}=1\,\mathrm{MHz}$.
@@
Instead of being a limit point that should be asymptotically approached, the completely vanishing field in cavity $1$ emerges at a midpoint of tuning the coupling $J_{2}$. As represented by the samples in Fig. 4(a), tuning the coupling strength from $J_{2}>J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ to $J_{2}<J_{2,c}$ will continuously switch the transparency window center from less than the $100\%$ transmission to more than the $100\%$ transmission, while the dispersion property demonstrating the slow-light effect is still preserved as seen from Fig. 4(b). As the distance between cavity $2$ and $3$ keeps increasing from where cavity $1$ is in the vacuum state of completely eliminated intracavity field, which is at $J_{2,c}=0.6$ MHz for the system in Fig. 4, more photons entering cavity $1$ realize an intracavity field phase to strengthen the output, so that a higher and higher ratio $|a_{out}/a_{in}|^{2}>1$ will be achieved until the system loses its dynamical stability at $J_{2}=0.2449$ MHz. Then it becomes something like two coupled resonators in a PT-symmetry broken regime due to the choice $J_{1}<\sqrt{\kappa_{1}\kappa_{2}}$. There is no restriction on the qualities of used optical resonators for seeing these phenomena of amplifying and reducing the output without modifying the built-in optical gain, since the system stability under an arbitrary net gain rate $\kappa_{2}$, which is detailed in Appendix D, and an available coupling strength range covering the point $J_{2,c}$ guarantee their occurrence. The structure in Fig. 1(a) can thus beat a common notion that the better the used optical resonators are, the better the induced transparency will be.